The taste of sour milk is lactic acid. What is the molecular formula for lactic acid if the percent composition is 40.00% C, 6.71% H, 53.29% O, and the approximate molar mass is 90 g/mol?
1
Expert's answer
2012-09-14T09:23:55-0400
90g/mol is 100%
for C --40% is 90*40/100=36 (36/12 is 3 C3)
for H = 6
(6/1=6 H6)
for O =48 (48/16= 3 O3)

formula is C3H6O3